OpenStudy (anonymous):

@amistre64 Three employees work at a shipping warehouse. Tom can fill an order in s minutes. Paco can fill an order S-2 in minutes. Carl can fill an order in s+1 minutes. When Tom and Paco work together, they take about 1 minute and 20 seconds to fill an order. When Paco and Carl work together, they take about 1 minute and 30 seconds to fill an order. a) How long does each person take to fill an order? b) How long would all three of them, working together, take to fill an order?

OpenStudy (anonymous):

So u do 1x30 and get the answer for a

OpenStudy (anonymous):

?

OpenStudy (anonymous):

what is 30x1

OpenStudy (anonymous):

30

OpenStudy (anonymous):

you get 30 ok

OpenStudy (anonymous):

lol that is not the answer...

OpenStudy (anonymous):

for a at least

OpenStudy (anonymous):

nope, how long does each employee take to fill one order. 30 is not the answer

OpenStudy (anonymous):

well then idk

OpenStudy (amistre64):

what equations can we form with the given information in the problem ...

OpenStudy (anonymous):

thats actually what i'm having trouble with, I don't know how to go about forming the equation

OpenStudy (anonymous):

well we know s is the time tom takes fill an order, and paco take 2 min less and carl takes 1 min more.

OpenStudy (amistre64):

we are given that t = this p = that c = another then we are given that working together: t+p = something and p+c = something else so its all a matter of substitution

OpenStudy (anonymous):

right.

OpenStudy (amistre64):

letting t=s ... t + (t-2) = 60+20 (t-2) + (t+1) = 60+30

OpenStudy (amistre64):

i dint read minutes, and converted to seconds ...

OpenStudy (amistre64):

60 + 20 seconds = 1 + 2/6 minutes

OpenStudy (anonymous):

okay. so should i do elimination?

OpenStudy (amistre64):

i don tthink its so much elimination anymore as simply solving for t

OpenStudy (anonymous):

I got t=41

OpenStudy (amistre64):

t + (t-2) min = 1+ 20/60 2t -2 = 1+ 20/60 2t = 3+ 20/60 t = 3/2+ 10/60 minutes

OpenStudy (anonymous):

that does not make sense... cuz then paco will have a negative time.

OpenStudy (amistre64):

hmm, is there an error in that?

OpenStudy (amistre64):

yeah, i noticed

OpenStudy (anonymous):

Someone else did this question, but i do not understand how they, did it, maybe you can explain it to me? https://ca.answers.yahoo.com/question/index?qid=20101017101534AA8UA9t

OpenStudy (amistre64):

it takes t and p 1,20 to finish a project it takes p and c 1,30 to finish a project by proportions t/1,20 + p/1,20 = 1 project assuming the same project p/1,30 + c/1,30 = 1 project somethings churning in there

OpenStudy (anonymous):

oh i c

OpenStudy (amistre64):

elimination would only be useful of we had 3 equations in 3 unknowns, but we have 2 equations in 3 unknowns

OpenStudy (anonymous):

I'm stumped on this question, hmm

OpenStudy (anonymous):

I dont understand why the person who solved this did this 1/s + 1/(s - 2) = 1/(4/3) = 3/4

OpenStudy (amistre64):

im not so sure why the did either ... if its accurate then its by some other method that they are more comfortable with

OpenStudy (amistre64):

s/1,20 + (s-2)/1,20 = 1 project (s-2)/1,30 + (s+1)/1,30 = 1 project s/1,20 + (s-2)/1,20 = (s-2)/1,30 + (s+1)/1,30 s/1,20 + s/1,20 -2/1,20 = s//1,30 -2/1,30 + s/1,30 +1/1,30 s/1,20 + s/1,20 -2/1,20 = s//1,30 -2/1,30 + s/1,30 +1/1,30

OpenStudy (anonymous):

Still does not make sense, i guess im just going to have to ask my teacher how to go about solving this question. Thanks for having a look at it.

OpenStudy (amistre64):

t=5, p=3, c=6 but not this seems odd to me as well. srry, i cant seem to think straight on this one

OpenStudy (amistre64):

t takes s minutes to do 1 job; the amount of the job per minutes is therefore: 1/s of the job each minute p take s-2 mintes to do 1 job, the amount of a job that p gets done in 1 minutes is thereofre 1/(s-2) c takes ... same concept this is better

OpenStudy (anonymous):

oh ic you put it into a rate.

OpenStudy (amistre64):

yes, finding a unit rate is important so that we can standardize their work efforts

OpenStudy (amistre64):

i wonder if calculus would help :) id prolly mess it up if i tried rates of change

OpenStudy (amistre64):

to do one job, \[t+p = 8/6~min \] so in 1 minute they can do 6/8 of the job the unit rate

OpenStudy (anonymous):

\[\frac{ 1 }{ s }+\frac{ 1 }{ s-2 }=\frac{ 4 }{ 3 }\] I think then 4/3 become 3/4 since what ever we do to one side we do to the other, correct me if im wrong.

OpenStudy (amistre64):

4/3 is not a unit rate, its a total time frame like s or s-1 we want a unit rate for comparison so 3/4 is what we are looking for i believe

OpenStudy (anonymous):

well then, now everything works out. and i got s=4

OpenStudy (anonymous):

so it take tom 4min, paco 2min, and carl 5min.

OpenStudy (amistre64):

then they can do 1/4 and 1/2 and 1/5 of a job per minute 1/4 + 1/2 = 3/4 of the job in 1 minute i like it

OpenStudy (anonymous):

Alright thanks for helping with this abomination of a question lol.

OpenStudy (amistre64):

4/3 (1/4 + 1/2) = 1 job (1/3 + 2/3) = 1 job in 4/3 minutes

OpenStudy (amistre64):

i swear this is simple enough to accomplish when you recall what to do lol